# Modeling sea ice in the marginal ice zone as a dense granular flow with rheology inferred from a discrete element model

### v1

@@ Appendix B Existence and uniqueness of solutions to the momentum equation
Since the momentum equation (4.8) of the continuous model is equivalent to the minimization of the functional $\mathcal{J}$ over the space $V$, defined in (4.5), it suffices to show that $\mathcal{J}$ admits a unique minimizer to prove the existence and uniqueness of solutions to the momentum equation. To do so, we first write the functional $\mathcal{J}:V\to\mathbb{R}$ as follows for simplicity:
@@
According to (Evans,, 2022, theorem 2, chapter 8), at least one function $V$ exists that minimizes $\mathcal{J}$ if the functional is convex and coercive. It is straightforward to check that $\mathcal{J}$ is convex. The functional $\mathcal{J}$ is said to be coercive in $V$ if
@@
Acknowledgements. We thank Andrew Thompson for many insightful comments and suggestions shared with us during the preparation of the manuscript.
Funding.This work has been supported by the Multidisciplinary University Research Initiatives (MURI) Program, Office of Naval Research (ONR) grant number N00014-19-1-242.
Author contributions. All authors contributed to conceiving and designing the study. GD wrote the paper, designed the figures, performed the simulations and mathematical analysis. MG, SG and GS revised the paper. MG, RH and SG contributed to the simulations, and SG contributed simulation tools. MG, GS and RH provided input to the analysis.

### v2

@@ Appendix B Existence and uniqueness of solutions to the momentum equation
Since the momentum equation (4.8) of the continuum model is equivalent to the minimization of the functional $\mathcal{J}$ over the space $V$, defined in (4.5), it suffices to show that $\mathcal{J}$ admits a unique minimizer to prove the existence and uniqueness of solutions to the momentum equation. To do so, we first write the functional $\mathcal{J}:V\to\mathbb{R}$ as follows for simplicity:
@@
According to (Evans, 2022, theorem 2, chapter 8), at least one function $V$ exists that minimizes $\mathcal{J}$ if the functional is convex and coercive. It is straightforward to check that $\mathcal{J}$ is convex. The functional $\mathcal{J}$ is said to be coercive in $V$ if
@@
[…]
[Acknowledgements]We thank Andrew Thompson for many insightful comments and suggestions shared with us during the preparation of the manuscript.
[…]
[Funding]This work has been supported by the Multidisciplinary University Research Initiatives (MURI) Program, Office of Naval Research (ONR) grant number N00014-19-1-242.
[…]
[Declaration of Interests]The authors report no conflict of interest.
[…]
[Author contributions]All authors contributed to conceiving and designing the study. GD wrote the paper, designed the figures, performed the simulations and mathematical analysis. MG, SG and GS revised the paper. MG, RH and SG contributed to the simulations, and SG contributed simulation tools. MG, GS and RH provided input to the analysis.